One mole of an ideal gas, cp = (7/2)R and Cv =(5/2)R is compressed adiabatically in a piston cylinder device from 2 bar and 25℃ to 7 bar. The process is irreversible and requires 35% more work than a reversible, adiabatic compression from the same initial state to the final pressure. What is entropy change of the gas?
1
Expert's answer
2020-02-27T09:02:53-0500

"\Delta"S = n(Cpln(T2/T1)-R*ln(p2/p1))

p2 = 7 bar

p1 = 2 bar

T1 = 25+273 = 298 K

If reversible:

W = "\alpha"*n*R*T1*((p2/p1)^(("\gamma"-1))/"\gamma" )-1)

"\gamma" =Cp/Cv = 7/5 = 1.4

"\alpha" =1/("\gamma" -1) = 2.5

W = 2.5 * 8.314 * 298 * (3.5^(0.4/1.4)-1) = 2666 J

If irreversible:

W = 2666 *1.35 = 3599 J

Cv(T2-T1) = W

T2 = W/Cv +T1 = 3599/(5*8.314/2) + 298 = 471 K

ΔS = Cpln(T2/T1)-R*ln(p2/p1) = (7*8.314/2)*ln(471/298)-8.314*ln(7/2) = 2.91 J/K